Left on together the, cold and hot water faucets of a certain bathtub take five minutes to fill the tub. If it takes the hot water faucet 15 minutes to fill the tub by it self, how long will it take the Cold-water faucet to fill the tub on its own?
Answer by Boreal(15235)

In one minute, the cold fills (1/x) of the tub and the hot (1/15), so together it is [(1/x)+(1/15)]5=1 full tub
this is (5/x)+(1/3)=1
LCD is 3x
15+x=3x
2x=15
x=7.5 minutes for cold water to fill on own.
in 5 minutes, 2/3 full by cold and 5/15=1/3 for warm.